Find s and r if (x + 3) is a factor of sx³ + rx² - 28x + 15 and has a remainder of -60 when divided by (x - 3).


1
Expert's answer
2021-03-02T06:21:44-0500
"s(-3)^3+r(-3)^2-28(-3)+15=0"

"-9s+3r+28+5=0"

"3s-r=11"


"s(3)^3+r(3)^2-28(3)+15=-60"

"9s+3r-28+25=0"

"3s+r=1"

"\begin{matrix}
 3s-r=11 \\
 3s+r=1
\end{matrix}"

"\begin{matrix}
 r=3s-11 \\
 6s=12
\end{matrix}"

"\begin{matrix}
 s=2 \\

 r=-5
\end{matrix}"